Observers: Linhua Jiang, Emily McLinden

7:20 pm: technical tests by Olga.

8:00 pm: pointing to Emily's target obj1.

8:20 pm: struggling with guide star, turned out that catalog was wrong, finally we got a right guide star.

9:00 pm: now we're observing science targets after struggling with grating orders. H+K grating was expected to work on the 1st order, but it was put on the 2nd order. We fixed it.

9:20 pm: keep observing, sky clear and seeing good (<0.8").

9:40 pm: seeing getting worse.

10:00 pm: seeing is 1.6" now.

10:10 pm: seeing jumped to 2"-3".

10:30 pm: seeing varying between 2"~3".

10:50 pm: seeing varying between 2"~4".

10:55 pm: stop observing science targets due to terrible seeing.

2:10 am: seeing below 2", we slew to Ran's backup target J1339. Finish a pair of 5 min exposures.

3:00 am: seeing improves to 1", we slew to SDF.

3:25 am: start to observe SDF.

3:45 am: first image came out very ugly (strong persistence), also probably not well positioned.

3:50 am: re-acquisition, after many tests, we still couldn't put slits in the right positions. we double checked the positions of ref. stars, and they're very accurate. Then Olga realized that there might be some problem with transformation between ra,dec and x,y. Luci is not well calibrated because they lost almost all technical nights due to bad weather.

-- XiaohuiFan - 17 Apr 2012
Topic revision: r7 - 18 Apr 2012, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback